OpenStudy (anonymous):

(4-z/ z-6) - (3z-5/6-z) simplify if possible

OpenStudy (anonymous):

(2z-1)/(z-6)

OpenStudy (dumbcow):

Factor out a negative in 2nd fraction -> -(z-6) = (6-z) (4-z)/(z-6) + (3z-5)/(z-6) = (2z-1)/(z-6)

OpenStudy (anonymous):

thanks